OpenStudy (anonymous):

welling to give a metal if help for 3 questions just reply

OpenStudy (anonymous):

what is it

OpenStudy (anonymous):

Parker earned $32 by selling 27 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold. 8 10 12 14

OpenStudy (anonymous):

Jackson ordered some clothes for school online. He bought 2 pairs of pants and 5 shirts. He spent a total of $138.45. Part 1: Write the equation for this word problem. Be sure to identify your variables. Part 2: If each pair of pants cost $14.25, solve the equation to find the cost of each shirt. Show your work.

OpenStudy (anonymous):

Graph the ordered pairs (0, 4) and (1, -1) and connect to form a line. Which quadrant contains no point for this linear function? Explain how you found your answer

OpenStudy (nikato):

Wait, let's do one problem at a time

OpenStudy (nikato):

So for the first. Can u set up two equations? Let a be the number of small glass and b be the number of large glass

OpenStudy (anonymous):

ok a=small b=large glass

OpenStudy (nikato):

Those aren't really equation. They just show what each variable represents.

OpenStudy (nikato):

I'll give u a hint. One of the equation will equal 27 and the other 32

OpenStudy (anonymous):

S+L = 27 1S + 1.50 = 32 is this right

OpenStudy (nikato):

Yes except I think u left available. Lets say 1.50b then

OpenStudy (nikato):

Can u solve this system of equation?

OpenStudy (anonymous):

I try would divide the 1.50 into the 27 or 32

OpenStudy (nikato):

No, is substitution familiar to u?

OpenStudy (nikato):

In the first equation , u can say s=27-l

OpenStudy (nikato):

Then substitute to that for s in the second equation

OpenStudy (anonymous):

I honestly don't know

OpenStudy (nikato):

D u see how I got S=27-L

OpenStudy (anonymous):

you got them from the glass

OpenStudy (nikato):

Yea, u had S+L=27 and i subtracted L on both sides, making it S=27-L

OpenStudy (anonymous):

ok

OpenStudy (nikato):

Now what do u get when u substitute that^ to the second equation 1S+1.5L=32

OpenStudy (anonymous):

you do the same to the other equation right

OpenStudy (nikato):

No were using substitution

OpenStudy (anonymous):

would it be like subtract 1.5 on both sides now

OpenStudy (nikato):

No just substitr for S

OpenStudy (nikato):

So 1(27-L)+1.5L=32 Do u see this? How this is done?

OpenStudy (anonymous):

yes now do you use the destitute property

OpenStudy (nikato):

Yea,

OpenStudy (nikato):

Can u try to solve for L

OpenStudy (anonymous):

don't l = 1

OpenStudy (nikato):

No

OpenStudy (nikato):

Distribute to get 27-1L+1.5L=32

OpenStudy (nikato):

Combine like terms 27+0.5L=32 Subtract 27 on both sides 0.5L=5 Then divide both sides by 0.5 and u will get L

OpenStudy (anonymous):

12

OpenStudy (nikato):

Try again 5/0.5

OpenStudy (anonymous):

10

OpenStudy (anonymous):

do u want me to give your metal now

OpenStudy (nikato):

Yes, and that would be ur answer becuz u are looking for the large glass and in the beginning we said that L= large glass

OpenStudy (nikato):

Oh, it's fine. Do u still still need help with ur other question? If u don't mind, can u start a new one and tag me becuz this is getting too long

OpenStudy (anonymous):

no I don't mind you can do one more and the third one I could try myself

OpenStudy (nikato):

Ok then. U can start a new one and tag me whenever u feel like it

OpenStudy (anonymous):

ok here is your medal thanks for helping